A box is being constructed by cutting 3-inch squares from the corners of a rectangular sheet of metal that is 4 inches longer than it is wide. If the box is to have a volume of 135 cubic inches, find the dimensions of the metal sheet.
Answer by ankor@dixie-net.com(22740)

A box is being constructed by cutting 3-inch squares from the corners of a
rectangular sheet of metal that is 4 inches longer than it is wide.
If the box is to have a volume of 135 cubic inches, find the dimensions of the metal sheet.
:
Assume this is an open box
Height = 3 inches
:
Let x = width of the box
then
(x+4) = length of the box
:
Volume
3x(x+4) = 135
3x^2 + 12x - 135 = 0
Factor
(3x + 27)(x - 5) = 0
positive solution
x = 5 inches is the width
then
5+4 = 9 inches is the length
:
Cutting the 3 inches corners adds 6 inches to each dimension
Metal sheet = 15 by 11 inches